# Where does the simplified Stellar Contamination Model fail in Exoplanet Transmission Spectroscopy?

### v1

@@
Transmission spectroscopy has become a cornerstone technique for probing exoplanetary atmospheres, enabling the detection of chemical species, clouds, and scattering phenomena (Seager.and.Sasselov.2000ApJ...537..916S; Brown.et.al.2001ApJ...552..699B; Charbonneau.et.al.2002ApJ...568..377C). However, interpretation of these observations can be made more challenging by the presence of unocculted stellar photospheric heterogeneities (i.e., starspots and faculae), which distort the host star’s spectrum during transits and can mimic, or obscure, planetary atmospheric features (McCullough.et.al.2014ApJ...791...55M; Herrero.et.al.2016A&A...586A.131H; Rackham.et.al.2017ApJ...834..151R; Rackham.et.al.2018ApJ...853..122R; Rackham.et.al.2019; Chachan.et.al.2019AJ....158..244C; Iyer.and.Line.2020ApJ...889...78I).
This source of systematic distortion, sometimes termed the Transit Light Source Effect (TLSE, Rackham.et.al.2018ApJ...853..122R), can produce apparent spectral modulation that surpasses the amplitude of genuine planetary atmospheric signals, particularly in systems with active M-dwarfs and K-type stars. Although the James Webb Space Telescope (JWST) has unprecedented sensitivity to molecules such as CH4, CO2, and H2O, wavelength-dependent flux deficits introduced by unocculted spots and faculae can mimic molecular absorption bands, mask genuine spectral features, and bias the retrieved planet-to-star radius ratio, thereby complicating the interpretation of transmission spectra.
Rackham.et.al.2018ApJ...853..122R introduced an analytic formalism – hereafter referred to as the Rackham-TLSE (R-TLSE) – which has since become the default means of mitigating stellar heterogeneity in atmospheric retrievals (e.g., Lim.et.al.2023ApJ...955L..22L; May.et.al.2023ApJ...959L...9M; Moran.et.al.2023ApJ...948L..11M; Cadieux.et.al.2024ApJ...970L...2C; Fournier-Tondreau.et.al.2024MNRAS.528.3354F; Petit.dit.de.la.Roche.et.al.2024A&A...692A..83P; Radica.et.al.2025ApJ...979L...5R; Saba.et.al.2025ApJS..276...70S; Perdelwitz.et.al.2025ApJ...980L..42P). Although the R-TLSE formalism has become a widely used approximation to model stellar contamination, it remains a simplified treatment of the effects of photospheric inhomogeneity. This approach describes the observed transit depth as a function of the contrast between the spectrum of the pristine photosphere (), the spectrum of heterogeneity (), and the fractional coverage of the stellar disc by active regions (the filling factor). However, it does not account for the geometry of the transit chord, the stellar limb-darkening profile, and the spatial distribution of active regions across the stellar disc. As a result, atmospheric retrievals based solely on the TLSE formalism may be biased, particularly for systems with complex stellar activity that are observed with high-precision measurements.
@@
The structure of this paper is as follows. In section 2, we describe in detail the implementation of the ECLIPSE-X model and the treatment of the Rackham–TLSE prescription. In section 3, we first present an idealised, noise-free comparison that employs a differential error metric to quantify discrepancies between the two formalisms for the three archetypal systems aforementioned. We then apply both frameworks to the JWST/NIRISS SOSS transmission spectrum of LHS 1140 b to assess their performance in a realistic observational context. We also discuss the implications of these findings for exoplanet atmosphere characterisation, focussing on the regimes in which the R–TLSE correction loses validity while noting the specific conditions where it remains a useful simplification. The main conclusions are summarised in section 4. Supplementary figures and retrieval outputs for K2-18 b and WASP-69 b are provided in Appendix A.
@@ 2.1 The ECLIPSE-X Starspot model
The ECLIPSE-X111https://github.com/ViktorSumida/ECLIPSE-Xlambda (ECLIPSE-Xlambda, v1.0.0, Zenodo, doi:10.5281/zenodo.10888850, as developed on GitHub) is an advanced transit modelling tool built upon the foundation of the ECLIPSE222https://github.com/Transit-Model-CRAAM/ model, originally proposed by Silva.2003. The code expanded the functionality of ECLIPSE allowing for simulation of transit light curves across a wide range of wavelengths. The original ECLIPSE code was primarily designed to model active-region occultations using a 2D image of a synthetic star with spots on the surface of the limb-darkened stellar disc. By analysing the photometric variations in the observed light curves during exoplanet transits, the physical properties of these systems can be inferred (e.g., Silva-Valio.et.al.2010; Valio.et.al.2017; Zaleski.et.al.2019; Zaleski.et.al.2020; zaleski.et.al.2022; Netto.and.Valio.2020; Araujo.and.Valio.2021; valio2022; Valio2024).
To produce synthetic light curves, ECLIPSE-X requires a 3D matrix to store stellar disc data at given wavelengths. This matrix encapsulates the spatial information of the star’s surface, enabling the code to simulate the effect of the transit and accurately compute the resulting flux variations. ECLIPSE-X generates multiwavelength synthetic light curves that closely resemble the observed data, and allows calculation of the transit depth which in turn enables the analysis of the atmospheric properties of the exoplanetary system under investigation. The code is primarily written in Python although functions in C were implemented to enhance computational efficiency. Next, we detail the relevant input parameters of ECLIPSE-X. Additional input parameters and their details can be found in Silva.2003.
@@
2.1.1 Constructing wavelength-dependent contrasts
In both the ECLIPSE-X and R–TLSE (see subsection 2.2) frameworks, stellar contamination can be expressed in terms of the wavelength-dependent contrast between the heterogeneous region and the pristine photosphere,
where and are the disc-integrated fluxes from the heterogeneous region and the immaculate photosphere, respectively. When is computed directly from high-resolution stellar-atmosphere spectra, for instance by interpolating PHOENIX specific intensities (Hauschildt.et.al.1997ApJ...483..390H; Hauschildt.et.al.1999JCoAM.109...41H; Husser.et.al.2013) with LDTK (Parviainen.and.Aigrain.2015MNRAS.453.3821P), it exhibits sharp, rapidly varying structure associated with atomic and molecular lines. Within a pixelated transit model that already includes wavelength-dependent limb darkening, feeding these line-dominated contrasts directly into the calculation leads to a strong and largely artificial amplification of individual spectral lines: the effective limb-darkening coefficients track these narrow features (see Figure 6) and imprint them onto the transit depth in a way that is not present in the original one-dimensional atmosphere models.
To prevent this line-amplification problem, in this work we decouple the smooth continuum behaviour of the stellar intensity from the detailed line structure. For the computation of limb-darkening coefficients, we do not use monochromatic PHOENIX intensities, as this would propagate line structure directly into the limb-darkening law. Instead, we adopt a plane-parallel grey-atmosphere approximation in radiative equilibrium to obtain a spectrally smooth centre-to-limb intensity profile. For each wavelength bin, we compute the emergent intensity as a function of , normalise by , and fit the four-parameter limb-darkening law of Claret.2000. The best-fitting coefficients are obtained by minimising the squared residuals between the grey-atmosphere intensity profile and the parametric law using a Levenberg–Marquardt non-linear least-squares algorithm, following the same optimisation strategy adopted by Claret.et.al.2025A&A...699A..97C for PHOENIX-based JWST passbands. This yields a set of limb-darkening coefficients that varies smoothly with wavelength and is free from narrow spectral-line signatures. The same grey-atmosphere limb-darkening law is used for both the immaculate photosphere and the active regions, while their relative brightness is set by the black-body contrast:
@@ 2.1.1 Constructing wavelength-dependent contrasts
For the self-consistent ECLIPSE-X simulations, the smooth component is then replaced by the black-body contrast implied by Equation 3, combined with the grey-atmosphere limb-darkening description aforementioned, while the residual is applied as a multiplicative, wavelength-dependent correction at the level of the disc-integrated fluxes. In this way, the contaminated signal predicted by ECLIPSE-X is governed by a spectrally smooth continuum and physically motivated limb darkening, while the fine spectral structure of realistic stellar spectra is still incorporated consistently into the model.
2.2 R-TLSE Model Comparision
According to Rackham.et.al.2018ApJ...853..122R, the parameter represents the correction factor introduced by the TLSE. It quantifies the discrepancy between the observed transit depth and the intrinsic transit depth expected in the absence of stellar heterogeneities at a given wavelength . The factor is defined as the ratio of the observed transit depth to the intrinsic transit depth, accounting for the presence of unocculted stellar spots and faculae. Mathematically, it is expressed as:
@@ 2.2 R-TLSE Model Comparision
In this study, we compare the simple R-TLSE model proposed by Rackham.et.al.2019 to our TLSE model, which incorporates additional physical effects. The comparison in this study will be carried out in a straightforward and systematic manner. First, we simulate two scenarios: a pristine, unspotted star (serving as a baseline) and a star with heterogeneous surface features (incorporating stellar contamination). For the TSLE, we will use the unspotted star as a reference and infer the stellar contamination through the contamination factor , allowing for a direct evaluation against the theoretical predictions of the R-TLSE model. To quantify discrepancies between the frameworks, we introduce the error:
where is the transit depth obtained from our simulations, and is the transit depth computed using Equation 5 of R-TSLE (Rackham.et.al.2018ApJ...853..122R). This metric isolates systematic biases introduced by the R-TLSE’s simplified treatment of limb darkening and transit geometry. Positive values of Error indicate that R–TLSE predicts a shallower transit than ECLIPSE-X (i.e., it overcorrects for stellar contamination), whereas negative values indicate that R–TLSE predicts a deeper transit and thus undercorrects for stellar contamination.
@@ 3.1 Idealised model–model comparison
Case 1 – Fixed filling factor, variable temperature contrast: The filling factor was set to 0.15 for both starspots and faculae, while the temperature contrast was varied from K to K. This disentangles spectral biases driven purely by temperature differences. The adopted upper limit for the filling factor () is motivated by recent atmospheric retrievals based on JWST transmission spectra, which inferred comparable starspot and faculae coverages in M-dwarf systems. Specifically, May.et.al.2023ApJ...959L...9M and Cadieux.et.al.2024ApJ...970L...2C reported coverages of up to 40% for GJ 1132 b and 30% for LHS 1140 b, based on NIRSpec and NIRISS data, respectively.
Case 2 – Fixed temperature contrast, variable filling factor: A constant temperature contrast of K was applied, while the filling factor was varied from 0.01 to 0.15. This isolates the effect of the spatial coverage of active regions. We constrained the temperature contrast between active regions and the photosphere to K, consistent with values inferred from those retrievals (May.et.al.2023ApJ...959L...9M; Cadieux.et.al.2024ApJ...970L...2C) and from solar analogue studies (e.g., Meunier.et.al.2010A&A...512A..39M).
@@
The pattern observed in the curves in Figures 1 and 2 – and Figure 7 in Appendix A – aligns with expectations: the largest wavelength-dependent errors between the R-TLSE approximation and the ECLIPSE-X model occur at wavelengths where limb darkening induces the most pronounced brightness gradients across the stellar disc, particularly in the optical regime (see Alexoudi.et.al.2020A&A...640A.134A). For LHS 1140 b, our simulations reveal significant discrepancies, especially at optical wavelengths. At wavelengths shorter than 800 nm, systematic biases exceed 50 ppm, reaching maxima of approximately 170 ppm for unocculted starspots and 140 ppm for faculae near 500 nm.
These offsets arise from the R-TLSE’s neglect of wavelength-dependent limb-darkening contrasts and the geometric modulation of stellar contamination. At longer wavelengths (1000–2000 nm), the error diminishes to below 10 ppm, as limb-darkening effects weaken and stellar spectral energy distributions become smoother. Adopting the system parameters of Cadieux.et.al.2024ApJ...970L...2C, a single atmospheric scale height of an N2-dominated atmosphere on LHS 1140 b corresponds to only a few ppm (on the order of 4–5 ppm) in transit depth; the 50–170 ppm optical discrepancies between ECLIPSE-X and the R–TLSE approximation therefore amount to on the order of 10–40 atmospheric scale heights. This wavelength-dependent bias, therefore, has direct implications for atmospheric retrievals and for the interpretation of slope-like features in terms of Rayleigh scattering or aerosols. Previous TLSE-based retrieval studies have already shown that specific stellar modelling assumptions and particular treatments of stellar contamination can imprint or distort such apparent slopes, biasing the inferred atmospheric properties (see, for example, Zhang.et.al.2018AJ....156..178Z; Iyer.and.Line.2020ApJ...889...78I; Rackham.and.deWit.2024AJ....168...82R).
In their study, Cadieux.et.al.2024ApJ...970L...2C report a marginal detection of N2 in the atmosphere of the temperate super-Earth LHS 1140 b at a significance level of 2.3. This result is primarily driven by a modest short-wavelength spectral slope (500–800 nm), which they interpret as Rayleigh scattering from a high-MMW atmosphere dominated by N2. Although the observed transmission spectrum remains largely flat at infrared wavelengths (1 m), it exhibits a subtle blueward slope in the visible regime – precisely where the R-TLSE approximation has the greatest potential for errors. The same analysis also reveals a statistically significant detection (5.8) of unocculted stellar faculae, covering approximately 20% of the stellar disc. This contamination is modelled using the R-TLSE formalism. However, despite the inclusion of this correction, the fit to the data exhibits a relatively high reduced chi-squared, hinting that the R-TLSE may not fully capture the spectral distortions induced by photospheric heterogeneities.
@@
For the sub-Neptune K2-18 b (see Figure 7, first and second rows), our simulations show that in the optical range (500–600 nm), the R-TLSE approximation and our self-consistent model can differ by about 80–100 ppm. Such offsets are comparable to the amplitude of key atmospheric signals in sub-Neptune spectra. Beyond 1.5 m, however, these discrepancies drop below 10 ppm, indicating that the R-TLSE remains reasonably accurate in the near-infrared, consistent with previous investigations that robustly detected methane (CH4) via R-TLSE-based equation (e.g., Madhusudhan.et.al.2023ApJ...956L..13M; Schmidt.et.al.2025arXiv250118477S). Schmidt.et.al.2025arXiv250118477S incorporated stellar contamination models ( at K cooler than the photosphere), showing that while these corrections minimally affect CH4 inferences, they are crucial to avoid false positives in trace species such as CO2.
In contrast, Madhusudhan.et.al.2023ApJ...956L..13M found no detectable stellar activity in M dwarfs hosting Hycean candidates. However, the apparently quiescent behaviour of M-type hosts such as K2-18 may be partly driven by orbital geometry: for transits probing mid-latitude chords (), our ECLIPSE-X simulations show that a spotless photosphere already yields nearly wavelength-independent transit depths, effectively removing most of the geometrical slope. In this “flat-geometry” regime, even modest chromatic stellar contamination or weak atmospheric gradients can produce shallow slopes that are comparable to the current observational uncertainties, making it difficult to distinguish genuinely high-MMW, cloud-dominated atmospheres from cases where residual stellar heterogeneity and geometry conspire to flatten the observed spectrum. This underscores the need for stellar-heterogeneity models that are explicitly consistent with the system’s orbital geometry. A more detailed discussion of the impact of limb darkening and transit geometry on transmission spectra is provided in Alexoudi.et.al.2020A&A...640A.134A and in Section 3.2.
@@
Such large discrepancies become even more significant in light of the recent VLT/FORS2 spectrophotometry of WASP-69 b by Petit.dit.de.la.Roche.et.al.2024A&A...692A..83P, covering 340–1100 nm. Their retrievals, which applied the R-TLSE correction equation to account for stellar contamination, inferred an extensive facula coverage of 30% with a temperature excess of K. For a K-type star, such a large facular filling factor is already somewhat extreme and is likely highly biased, given the simplified assumptions inherent to the correction method. Moreover, the planet’s mid-latitude impact parameter () implies that the transit chord probes intermediate stellar latitudes, close to the 40–45∘ regime where our ECLIPSE-X simulations predict nearly wavelength-independent transit depths for a spotless photosphere. In this resulting “flat-geometry” configuration, most of the purely geometrical slope is removed, so the negative optical slope seen in the FORS2 transmission spectrum is expected to arise primarily from the combined effects of stellar heterogeneity and atmospheric gradients, rather than from geometry alone.
@@
Additionally, a large facula filling factor naturally induces a negative spectral slope in the optical, as illustrated in Figure 3 of Petit.dit.de.la.Roche.et.al.2024A&A...692A..83P. In our self-consistent ECLIPSE-X calculations, the disagreement with the R-TLSE-based predictions grows systematically with increasing facula coverage: for high filling factors, the R-TLSE prescription produces much stronger chromatic stellar contamination than the pixel-resolved model. This behaviour suggests that, when applied in the mid-latitude geometric regime of WASP-69 b, the R-TLSE approach can substantially overpredict the level of stellar contamination and thus overestimate the facula filling factor.
@@
Prior atmospheric analyses based on HST/STIS and WFC3 data suggested strong Rayleigh scattering slopes, which were interpreted as evidence for high-altitude aerosols (e.g., Estrela.et.al.2021AJ....162...91E). However, the steep slope reported in that study might instead result from unocculted starspots or faculae, since our contamination-free simulations yield a flat spectral baseline for WASP-69 b. Therefore, it remains uncertain whether these optical slopes arise from genuine atmospheric properties or are artifacts of stellar heterogeneity.
@@
Meanwhile, Murgas.et.al.2020A&A...641A.158M reported an even steeper optical slope with GTC/OSIRIS, consistent with Rayleigh scattering, which they modelled with and ( K, K). Such extreme filling factors almost certainly do not trace stellar heterogeneity alone, but rather reflect a degeneracy in which part of the atmospheric signal is absorbed into the stellar-contamination term by the R-TLSE formalism. For instance, a spot filling factor of corresponds to an equivalent circular heterogeneous region with radius ; when the facular contribution is included, the total coverage () implies . In our 2D stellar model, such a vast coherent active region on the visible hemisphere would be difficult to reconcile with the mid-latitude transit chord of WASP-69 b without the planet occulting a substantial fraction of these inhomogeneities. This strongly suggests that the steep OSIRIS slope is more naturally explained by a combination of genuine atmospheric opacity and more moderate stellar heterogeneity than by the extreme active-region coverages implied by the disc-averaged R-TLSE prescription.

### v2

@@
Transmission spectroscopy has become a cornerstone technique for probing exoplanetary atmospheres, enabling the detection of chemical species, clouds, and scattering phenomena (Seager and Sasselov, 2000; Brown et al., 2001; Charbonneau et al., 2002). However, interpretation of these observations can be made more challenging by the presence of unocculted stellar photospheric heterogeneities (i.e., starspots and faculae), which distort the host star’s spectrum during transits and can mimic, or obscure, planetary atmospheric features (McCullough et al., 2014; Herrero et al., 2016; Rackham et al., 2017, 2018, 2019; Chachan et al., 2019; Iyer and Line, 2020).
This source of systematic distortion, sometimes termed the Transit Light Source Effect (TLSE, Rackham et al., 2018), can produce apparent spectral modulation that surpasses the amplitude of genuine planetary atmospheric signals, particularly in systems with active M-dwarfs and K-type stars. Although the James Webb Space Telescope (JWST) has unprecedented sensitivity to molecules such as CH4, CO2, and H2O, wavelength-dependent flux deficits introduced by unocculted spots and faculae can mimic molecular absorption bands, mask genuine spectral features, and bias the retrieved planet-to-star radius ratio, thereby complicating the interpretation of transmission spectra.
Rackham et al. (2018) introduced an analytic formalism – hereafter referred to as the Rackham-TLSE (R-TLSE) – which has since become the default means of mitigating stellar heterogeneity in atmospheric retrievals (e.g., Lim et al., 2023; May et al., 2023; Moran et al., 2023; Cadieux et al., 2024; Fournier-Tondreau et al., 2024; Petit dit de la Roche et al., 2024; Radica et al., 2025; Saba et al., 2025; Perdelwitz et al., 2025). Although the R-TLSE formalism has become a widely used approximation to model stellar contamination, it remains a simplified treatment of the effects of photospheric inhomogeneity. This approach describes the observed transit depth as a function of the contrast between the spectrum of the pristine photosphere (), the spectrum of heterogeneity (), and the fractional coverage of the stellar disc by active regions (the filling factor). However, it does not account for the geometry of the transit chord, the stellar limb-darkening profile, and the spatial distribution of active regions across the stellar disc. As a result, atmospheric retrievals based solely on the TLSE formalism may be biased, particularly for systems with complex stellar activity that are observed with high-precision measurements.
@@
The structure of this paper is as follows. In Section 2, we describe in detail the implementation of the ECLIPSE-X model and the treatment of the Rackham–TLSE prescription. In Section 3, we first present an idealised, noise-free comparison that employs a differential error metric to quantify discrepancies between the two formalisms for the three archetypal systems aforementioned. We then apply both frameworks to the JWST/NIRISS SOSS transmission spectrum of LHS 1140 b to assess their performance in a realistic observational context. We also discuss the implications of these findings for exoplanet atmosphere characterisation, focussing on the regimes in which the R–TLSE correction loses validity while noting the specific conditions where it remains a useful simplification. The main conclusions are summarised in Section 4. Supplementary figures and retrieval outputs for K2-18 b and WASP-69 b are provided in Section A.
@@ 2.1 The ECLIPSE-X starspot model
The ECLIPSE-X111https://github.com/ViktorSumida/ECLIPSE-Xlambda (Sumida and Valio, 2024, v1.0.0, Zenodo, doi:10.5281/zenodo.10888850, as developed on GitHub) is an advanced transit modelling tool built upon the foundation of the ECLIPSE222https://github.com/Transit-Model-CRAAM/ model, originally proposed by Silva (2003). The code expanded the functionality of ECLIPSE allowing for simulation of transit light curves across a wide range of wavelengths. The original ECLIPSE code was primarily designed to model active-region occultations using a 2D image of a synthetic star with spots on the surface of the limb-darkened stellar disc. By analysing the photometric variations in the observed light curves during exoplanet transits, the physical properties of these systems can be inferred (e.g., Silva-Valio et al., 2010; Valio et al., 2017; Zaleski et al., 2019, 2020, 2022; Netto and Valio, 2020; Araújo and Valio, 2021; Valio and Araújo, 2022; Valio et al., 2024).
To produce synthetic light curves, ECLIPSE-X requires a 3D matrix to store stellar disc data at given wavelengths. This matrix encapsulates the spatial information of the star’s surface, enabling the code to simulate the effect of the transit and accurately compute the resulting flux variations. ECLIPSE-X generates multiwavelength synthetic light curves that closely resemble the observed data, and allows calculation of the transit depth which in turn enables the analysis of the atmospheric properties of the exoplanetary system under investigation. The code is primarily written in Python although functions in C were implemented to enhance computational efficiency. Next, we detail the relevant input parameters of ECLIPSE-X. Additional input parameters and their details can be found in Silva (2003).
@@
2.2 Constructing wavelength-dependent contrasts
In both the ECLIPSE-X and R–TLSE (see Section 2.3) frameworks, stellar contamination can be expressed in terms of the wavelength-dependent contrast between the heterogeneous region and the pristine photosphere,
where and are the disc-integrated fluxes from the heterogeneous region and the immaculate photosphere, respectively. When is computed directly from high-resolution stellar-atmosphere spectra, for instance by interpolating PHOENIX specific intensities (Hauschildt et al., 1997; Hauschildt and Baron, 1999; Husser et al., 2013) with LDTK (Parviainen and Aigrain, 2015), it exhibits sharp, rapidly varying structure associated with atomic and molecular lines. Within a pixelated transit model that already includes wavelength-dependent limb darkening, feeding these line-dominated contrasts directly into the calculation leads to a strong and largely artificial amplification of individual spectral lines: the effective limb-darkening coefficients track these narrow features (see Figure 6) and imprint them onto the transit depth in a way that is not present in the original one-dimensional atmosphere models.
To prevent this line-amplification problem, in this work we decouple the smooth continuum behaviour of the stellar intensity from the detailed line structure. For the computation of limb-darkening coefficients, we do not use monochromatic PHOENIX intensities, as this would propagate line structure directly into the limb-darkening law. Instead, we adopt a plane-parallel grey-atmosphere approximation in radiative equilibrium to obtain a spectrally smooth centre-to-limb intensity profile. For each wavelength bin, we compute the emergent intensity as a function of , normalise by , and fit the four-parameter limb-darkening law of Claret (2000). The best-fitting coefficients are obtained by minimising the squared residuals between the grey-atmosphere intensity profile and the parametric law using a Levenberg–Marquardt non-linear least-squares algorithm, following the same optimisation strategy adopted by Claret et al. (2025) for PHOENIX-based JWST passbands. This yields a set of limb-darkening coefficients that varies smoothly with wavelength and is free from narrow spectral-line signatures. The same grey-atmosphere limb-darkening law is used for both the immaculate photosphere and the active regions, while their relative brightness is set by the black-body contrast:
@@ 2.2 Constructing wavelength-dependent contrasts
For the self-consistent ECLIPSE-X simulations, the smooth component is then replaced by the black-body contrast implied by Equation 3, combined with the aforementioned grey-atmosphere limb-darkening description, while the residual is applied as a multiplicative, wavelength-dependent correction at the level of the disc-integrated fluxes. In this way, the contaminated signal predicted by ECLIPSE-X is governed by a spectrally smooth continuum and physically motivated limb darkening, while the fine spectral structure of realistic stellar spectra is still consistently incorporated into the model.
2.3 R-TLSE model comparison
According to Rackham et al. (2018), the parameter represents the correction factor introduced by the TLSE. It quantifies the discrepancy between the observed transit depth and the intrinsic transit depth expected in the absence of stellar heterogeneities at a given wavelength . The factor is defined as the ratio of the observed transit depth to the intrinsic transit depth, accounting for the presence of unocculted stellar spots and faculae. Mathematically, it is expressed as:
@@ 2.3 R-TLSE model comparison
In this study, we compare the simple R-TLSE model proposed by Rackham et al. (2019) to our TLSE model, which incorporates additional physical effects. The comparison in this study will be carried out in a straightforward and systematic manner. First, we simulate two scenarios: a pristine, unspotted star (serving as a baseline) and a star with heterogeneous surface features (incorporating stellar contamination). For the TSLE, we will use the unspotted star as a reference and infer the stellar contamination through the contamination factor , allowing for a direct evaluation against the theoretical predictions of the R-TLSE model. To quantify discrepancies between the frameworks, we introduce the error:
where is the transit depth obtained from our simulations, and is the transit depth computed using Equation 5 of R-TSLE (Rackham et al., 2018). This metric isolates systematic biases introduced by the R-TLSE’s simplified treatment of limb darkening and transit geometry. Positive values of Error indicate that R–TLSE predicts a shallower transit than ECLIPSE-X (i.e., it overcorrects for stellar contamination), whereas negative values indicate that R–TLSE predicts a deeper transit and thus undercorrects for stellar contamination.
@@ 3.1 Idealised model–model comparison
Case 1 – Fixed filling factor, variable temperature contrast: The filling factor was set to 0.15 for both starspots and faculae, while the temperature contrast was varied from K to K. This disentangles spectral biases driven purely by temperature differences. The adopted upper limit for the filling factor () is motivated by recent atmospheric retrievals based on JWST transmission spectra, which inferred comparable starspot and faculae coverages in M-dwarf systems. Specifically, May et al. (2023) and Cadieux et al. (2024) reported coverages of up to 40% for GJ 1132 b and 30% for LHS 1140 b, based on NIRSpec and NIRISS data, respectively.
Case 2 – Fixed temperature contrast, variable filling factor: A constant temperature contrast of K was applied, while the filling factor was varied from 0.01 to 0.15. This isolates the effect of the spatial coverage of active regions. We constrained the temperature contrast between active regions and the photosphere to K, consistent with values inferred from those retrievals (May et al., 2023; Cadieux et al., 2024) and from solar analogue studies (e.g., Meunier et al., 2010).
@@
The pattern observed in the curves in Figures 1 and 2 – and Figure 7 in Appendix A – aligns with expectations: the largest wavelength-dependent errors between the R-TLSE approximation and the ECLIPSE-X model occur at wavelengths where limb darkening induces the most pronounced brightness gradients across the stellar disc, particularly in the optical regime (see Alexoudi et al., 2020). For LHS 1140 b, our simulations reveal significant discrepancies, especially at optical wavelengths. At wavelengths shorter than 800 nm, systematic biases exceed 50 ppm, reaching maxima of approximately 170 ppm for unocculted starspots and 140 ppm for faculae near 500 nm.
These offsets arise from the R-TLSE’s neglect of wavelength-dependent limb-darkening contrasts and the geometric modulation of stellar contamination. At longer wavelengths (1000–2000 nm), the error diminishes to below 10 ppm, as limb-darkening effects weaken and stellar spectral energy distributions become smoother. Adopting the system parameters of Cadieux et al. (2024), a single atmospheric scale height of an N2-dominated atmosphere on LHS 1140 b corresponds to only a few ppm (on the order of 4–5 ppm) in transit depth; the 50–170 ppm optical discrepancies between ECLIPSE-X and the R–TLSE approximation therefore amount to on the order of 10–40 atmospheric scale heights. This wavelength-dependent bias, therefore, has direct implications for atmospheric retrievals and for the interpretation of slope-like features in terms of Rayleigh scattering or aerosols. Previous TLSE-based retrieval studies have already shown that specific stellar modelling assumptions and particular treatments of stellar contamination can imprint or distort such apparent slopes, biasing the inferred atmospheric properties (see, for example, Zhang et al., 2018; Iyer and Line, 2020; Rackham and de Wit, 2024).
In their study, Cadieux et al. (2024) report a marginal detection of N2 in the atmosphere of the temperate super-Earth LHS 1140 b at a significance level of 2.3. This result is primarily driven by a modest short-wavelength spectral slope (500–800 nm), which they interpret as Rayleigh scattering from a high-MMW atmosphere dominated by N2. Although the observed transmission spectrum remains largely flat at infrared wavelengths (1 m), it exhibits a subtle blueward slope in the visible regime – precisely where the R-TLSE approximation has the greatest potential for errors. The same analysis also reveals a statistically significant detection (5.8) of unocculted stellar faculae, covering approximately 20% of the stellar disc. This contamination is modelled using the R-TLSE formalism. However, despite the inclusion of this correction, the fit to the data exhibits a relatively high reduced chi-squared, hinting that the R-TLSE may not fully capture the spectral distortions induced by photospheric heterogeneities.
@@
For the sub-Neptune K2-18 b (see Figure 7, first and second rows), our simulations show that in the optical range (500–600 nm), the R-TLSE approximation and our self-consistent model can differ by about 80–100 ppm. Such offsets are comparable to the amplitude of key atmospheric signals in sub-Neptune spectra. Beyond 1.5 m, however, these discrepancies drop below 10 ppm, indicating that the R-TLSE remains reasonably accurate in the near-infrared, consistent with previous investigations that robustly detected methane (CH4) via R-TLSE-based equation (e.g., Madhusudhan et al., 2023; Schmidt et al., 2025). Schmidt et al. (2025) incorporated stellar contamination models ( at K cooler than the photosphere), showing that while these corrections minimally affect CH4 inferences, they are crucial to avoid false positives in trace species such as CO2.
In contrast, Madhusudhan et al. (2023) found no detectable stellar activity in M dwarfs hosting Hycean candidates. However, the apparently quiescent behaviour of M-type hosts such as K2-18 may be partly driven by orbital geometry: for transits probing mid-latitude chords (), our ECLIPSE-X simulations show that a spotless photosphere already yields nearly wavelength-independent transit depths, effectively removing most of the geometrical slope. In this “flat-geometry” regime, even modest chromatic stellar contamination or weak atmospheric gradients can produce shallow slopes that are comparable to the current observational uncertainties, making it difficult to distinguish genuinely high-MMW, cloud-dominated atmospheres from cases where residual stellar heterogeneity and geometry conspire to flatten the observed spectrum. This underscores the need for stellar-heterogeneity models that are explicitly consistent with the system’s orbital geometry. A more detailed discussion of the impact of limb darkening and transit geometry on transmission spectra is provided in Alexoudi et al. (2020) and in Section 3.2.
@@
Such large discrepancies become even more significant in light of the recent VLT/FORS2 spectrophotometry of WASP-69 b by Petit dit de la Roche et al. (2024), covering 340–1100 nm. Their retrievals, which applied the R-TLSE correction equation to account for stellar contamination, inferred an extensive facula coverage of 30% with a temperature excess of K. For a K-type star, such a large facular filling factor is already somewhat extreme and is likely highly biased, given the simplified assumptions inherent to the correction method. Moreover, the planet’s mid-latitude impact parameter () implies that the transit chord probes intermediate stellar latitudes, close to the 40–45∘ regime where our ECLIPSE-X simulations predict nearly wavelength-independent transit depths for a spotless photosphere. In this resulting “flat-geometry” configuration, most of the purely geometrical slope is removed, so the negative optical slope seen in the FORS2 transmission spectrum is expected to arise primarily from the combined effects of stellar heterogeneity and atmospheric gradients, rather than from geometry alone.
@@
Additionally, a large facula filling factor naturally induces a negative spectral slope in the optical, as illustrated in Figure 3 of Petit dit de la Roche et al. (2024). In our self-consistent ECLIPSE-X calculations, the disagreement with the R-TLSE-based predictions grows systematically with increasing facula coverage: for high filling factors, the R-TLSE prescription produces much stronger chromatic stellar contamination than the pixel-resolved model. This behaviour suggests that, when applied in the mid-latitude geometric regime of WASP-69 b, the R-TLSE approach can substantially overpredict the level of stellar contamination and thus overestimate the facula filling factor.
@@
Prior atmospheric analyses based on HST/STIS and WFC3 data suggested strong Rayleigh scattering slopes, which were interpreted as evidence for high-altitude aerosols (e.g., Estrela et al., 2021). However, the steep slope reported in that study might instead result from unocculted starspots or faculae, since our contamination-free simulations yield a flat spectral baseline for WASP-69 b. Therefore, it remains uncertain whether these optical slopes arise from genuine atmospheric properties or are artifacts of stellar heterogeneity.
@@
Meanwhile, Murgas et al. (2020) reported an even steeper optical slope with GTC/OSIRIS, consistent with Rayleigh scattering, which they modelled with and ( K, K). Such extreme filling factors almost certainly do not trace stellar heterogeneity alone, but rather reflect a degeneracy in which part of the atmospheric signal is absorbed into the stellar-contamination term by the R-TLSE formalism. For instance, a spot filling factor of corresponds to an equivalent circular heterogeneous region with radius ; when the facular contribution is included, the total coverage () implies . In our 2D stellar model, such a vast coherent active region on the visible hemisphere would be difficult to reconcile with the mid-latitude transit chord of WASP-69 b without the planet occulting a substantial fraction of these inhomogeneities. This strongly suggests that the steep OSIRIS slope is more naturally explained by a combination of genuine atmospheric opacity and more moderate stellar heterogeneity than by the extreme active-region coverages implied by the disc-averaged R-TLSE prescription.